# A Comparative Study of Conventional and Tripolar EEG for High-Performance Reach-to-Grasp BCI Systems

### v1

@@ 2.5 Feature Extraction and Classification
In our previous investigations, comparisons between EEG and tEEG, specifically in terms of SNR and ERPs, have consistently demonstrated the superiority of tEEG. Building upon these findings, the present study aims to further contrast the capabilities of EEG and tEEG in a real-world classification task, thereby bridging the gap between analytical comparisons and practical applications.
The primary objective during the feature extraction phase was to capture relevant information from the EEG and tEEG data. The time window [-1, 1] s was defined as the time region of interest (tROI), with 0 s corresponding to the movement onset. Wavelets, specifically the complex Morlet wavelet (cmor), were employed to decompose these segments into various frequency bands: delta, theta, alpha, beta, and gamma. For each of the identified frequency bands in every channel, four statistical features were extracted from the magnitude of the wavelet coefficients. These features are: Mean of the magnitude, Variance of the magnitude, Skewness of the magnitude, Kurtosis of the magnitude.
Given that there were 4 channels, each channel containing data from 5 frequency bands, and each band producing 4 features, the total feature count amounted to 80 for every data instance. With the features in hand, the subsequent step was classification. We employed four distinct machine learning algorithms: Support Vector Machine (SVM), Random Forest (RF), Extreme Gradient Boosting (XGBoost), and Linear Discriminant Analysis (LDA). These techniques were used for both multiclass and binary classification scenarios, covering a range of hand-grasping situations. To assess the effectiveness of these models, we implemented a 5-fold cross-validation method. This approach helped us gain a reliable measure of the models’ ability to generalize, as it calculated the average performance indicators across all the validation folds.
@@ 3.1 Comparative Analysis
The comparative analysis of SNRs between the tEEG and conventional EEG technologies revealed a consistent pattern favoring the tEEG method across various channels. As delineated in Figure 3, the grand average SNR values for all subjects were assessed at channels P3, C3, C4, and F5. Notably, the tEEG (green boxes) consistently exhibited higher median SNRs across all channels when compared to the conventional EEG (blue boxes). Specifically, the median SNR for the tEEG ranged from 56.24 dB at channel C3 to 59.56 dB at channel P3, which was the highest observed across all channels and technologies. In comparison, the conventional EEG technology presented lower median SNRs, with the lowest at channel C4 (53.92 dB) and the highest at channel P3 (54.57 dB). In addition to the superior median SNRs, the tEEG also demonstrated lower variability in SNR measurements among subjects, as evidenced by the narrower interquartile ranges in the green boxes. This reduced variance suggests that the tEEG provides more consistent and reliable SNR values, which is a critical aspect of EEG signal quality.
@@
Figure 4 presents the grand average functional connectivity between EEG channels, quantified across all subjects for both EEG and tEEG technologies. The heatmaps visually compare the connectivity strengths, with conventional EEG displayed on the left and tEEG on the right. The conventional EEG heatmap shows considerable functional connectivity, with several channel pairs exhibiting strong correlations, such as P3-F5 (0.82) and C3-F5 (0.8), suggesting significant shared signal components across these channels. This is indicative of a lower spatial resolution, where the activity from different brain regions may be detected simultaneously by multiple channels. Conversely, the tEEG heatmap illustrates substantially lower connectivity values for all channel pairs when compared to conventional EEG. This is particularly evident in the markedly reduced correlations for channel pairs P3-F5 (0.01) and C3-F5 (0.048). The overall lower connectivity values in the tEEG heatmap underscore its higher spatial resolution, demonstrating its ability to more effectively discriminate between the electrical activities of different brain areas.
Figure 5 contrasts the ERPs and Global Field Power (GFP) between EEG and tEEG for subject 1 as an example. The first row represents data from conventional EEG, while the second row displays results from tEEG. The conditions compared across columns include No Movement, Power Grasp, and Precision Grasp, with time zero marking the onset of movement. During the No Movement condition, both technologies demonstrate baseline neural activity with flat ERP traces and low GFP fluctuations, as anticipated when no motor action is performed. This condition establishes a comparative rest state for subsequent analyses of neural dynamics. The Power Grasp condition reveals a stark contrast between the EEG technologies at the moment of movement onset. Conventional EEG captures a clear peak in the ERP waveform, with a corresponding GFP elevation indicative of neural engagement. However, tEEG exhibits a more sharply defined ERP peak and a crisper GFP response, suggesting its enhanced capability to detect the neural activation associated with the grasping motion. For the Precision Grasp, the distinction between EEG and tEEG is even more pronounced. EEG captures the ERP response adequately, but tEEG delineates more intricate fluctuations in the ERP waveforms across channels, as shown by the varied trace colors. The GFP for tEEG is more complex, particularly around the movement onset, highlighting its potential to resolve finer neural responses related to the intricate motor task of precision grasping.
@@
Figure 6 showcases the wavelet time-frequency representations for conventional EEG and tEEG, analyzing the neural dynamics from one second before to one second after the onset of movement, which is marked by time zero on the x-axis. During the Power Grasp and Precision Grasp conditions, the time-frequency maps exhibit a high increase in activity, particularly noticeable around 500ms post-movement onset. The tEEG maps display a more pronounced increase in activity across a range of frequencies when compared to conventional EEG. This increased activity, especially evident in the beta frequency band, is indicative of the sensorimotor processing associated with the execution of the grasping tasks. In the No Movement condition, conventional EEG exhibits minimal activity, aligning with expectations of neural quiescence. However, tEEG reveals subtle low-frequency activities even in this resting state, indicating its heightened sensitivity to neural oscillations that conventional EEG may not capture. The observed differences in the time-frequency domain between EEG and tEEG underscore the potential of tEEG to provide enhanced temporal and spectral resolution, which could translate to more accurate identification of distinct neural patterns associated with various motor tasks.
@@ 3.2 Feature Extraction and Classification
Table 1 presents the binary classification accuracies for distinguishing between No-movement and Power Grasp tasks across different subjects using various machine learning models with accuracies reported for both conventional EEG and tEEG. Across all subjects, tEEG consistently outperformed conventional EEG in classification accuracy. The mean accuracy improvements with tEEG ranged from 3.76% in SVM to a notable 16.30% in LDA, indicating a substantial enhancement in the machine learning models’ ability to utilize the data captured by tEEG. The Random Forest classifier achieved the highest mean accuracy for tEEG at 91.53%, which underscores the effectiveness of this model in conjunction with the high-resolution data provided by tEEG. Similarly, LDA showed the least improvement, yet still produced a mean accuracy increase of 16.30% with tEEG. Individual subject analysis revealed that Subject 7 (S7) exhibited the highest classification accuracy with tEEG at 98.33% using the XGBoost model, while the lowest was for Subject 10 (S10) at 67.5% using LDA.
| Subjects | Random For. (%) | SVM (%) | XGBoost (%) | LDA (%) | ||||
| EEG | tEEG | EEG | tEEG | EEG | tEEG | EEG | tEEG | |
| S1 | 70 | 90 | 75 | 75.56 | 73.33 | 88.89 | 51.67 | 78.89 |
| S2 | 82.82 | 87.42 | 76.54 | 85.42 | 84.36 | 84.32 | 56.28 | 72.89 |
| S3 | 71.76 | 92.36 | 74.05 | 83.3 | 76.34 | 90.17 | 50.26 | 84.1 |
| S4 | 95 | 96.11 | 89.17 | 85.56 | 90.83 | 97.22 | 68.33 | 76.67 |
| S5 | 69.58 | 90.56 | 69.17 | 68.06 | 67.08 | 86.11 | 60.42 | 69.44 |
| S6 | 80 | 91.33 | 73 | 78.67 | 76 | 88.67 | 43 | 78.67 |
| S7 | 94.38 | 98.24 | 93.12 | 96.67 | 92.5 | 98.33 | 84.38 | 89.58 |
| S8 | 73.75 | 89.58 | 70 | 70 | 71.25 | 90 | 66.88 | 73.75 |
| S9 | 73.12 | 91.67 | 67.5 | 72.92 | 65.62 | 90 | 59.38 | 72.08 |
| S10 | 68.12 | 87.08 | 66.88 | 75.83 | 68.12 | 82.5 | 60 | 67.5 |
| Mean | 77.85 | 91.53 | 75.44 | 79.20 | 76.54 | 89.62 | 60.06 | 76.36 |
@@
Table 2 details the binary classification accuracies for No-movement versus Precision Grasp tasks, utilizing a range of machine learning models and comparing the performance between conventional EEG and tEEG. The data illustrate a consistent trend where tEEG provides a marked improvement in classification accuracy over conventional EEG. The mean accuracies are notably higher with tEEG, exhibiting an increase of 12.98% for Random Forest, 8.84% for SVM, 15.65% for XGBoost, and 15.21% for LDA. The XGBoost classifier shows the most substantial average increase in accuracy when using tEEG, highlighting the synergy between this model and the high-resolution data from tEEG. Notably, every subject experienced an accuracy improvement with tEEG, affirming its universal enhancement of classification performance. Subject 7’s results stand out with an exceptional 99.17% accuracy using the Random Forest classifier in conjunction with tEEG, while the lowest accuracy with this technology is seen in Subject 8, with 67.50% using LDA. This spread suggests that while tEEG generally improves classification outcomes, the extent of its impact can vary among individuals.
| Subjects | Random For. (%) | SVM (%) | XGBoost (%) | LDA (%) | ||||
| EEG | tEEG | EEG | tEEG | EEG | tEEG | EEG | tEEG | |
| S1 | 76.67 | 90 | 78.33 | 85.56 | 78.33 | 90 | 51.67 | 71.11 |
| S2 | 78.08 | 88.53 | 62.56 | 82.21 | 67.18 | 84.37 | 53.21 | 76.11 |
| S3 | 75.03 | 93.22 | 62.35 | 82.56 | 66.99 | 93.19 | 48.95 | 71.2 |
| S4 | 93.33 | 92.78 | 88.33 | 82.22 | 91.67 | 92.22 | 64.17 | 77.78 |
| S5 | 70.42 | 88.89 | 72.08 | 76.94 | 69.17 | 84.72 | 63.33 | 69.72 |
| S6 | 80 | 88.67 | 67 | 87.33 | 75 | 90.67 | 62 | 76 |
| S7 | 93.75 | 99.17 | 92.5 | 94.58 | 93.12 | 97.92 | 80.62 | 92.92 |
| S8 | 68.12 | 89.58 | 70.62 | 73.33 | 64.38 | 88.75 | 54.37 | 67.50 |
| S9 | 70.62 | 89.17 | 68.12 | 77.08 | 67.5 | 88.75 | 65 | 70.42 |
| S10 | 71.25 | 87.08 | 69.38 | 77.92 | 67.5 | 86.67 | 54.37 | 77.08 |
| Mean | 77.73 | 90.71 | 73.13 | 81.97 | 74.08 | 89.73 | 59.77 | 74.98 |
@@
Table 3 summarizes the binary classification accuracies for discriminating between Power Grasp and Precision Grasp tasks, utilizing various machine learning models for both conventional EEG and tEEG. A key observation from the table is that while tEEG demonstrates an overall improvement in classification accuracy for most models, the SVM model with conventional EEG data exhibits a slightly higher average accuracy (75.53%) compared to tEEG (74.23%). However, this trend does not hold across other models. The XGBoost classifier, in particular, shows a marked increase in mean accuracy from 72.59% with EEG to 86.59% with tEEG. This substantial improvement highlights the potential for tEEG to enhance classification performance when paired with models that can exploit its higher-resolution data. Subject 7’s near-perfect classification with tEEG using Random Forest underscores this potential. For Random Forest and LDA models, tEEG also outperforms conventional EEG, with mean accuracies higher by 13.28% and 12.41%, respectively. These results illustrate the capacity of tEEG to provide more discernible patterns of neural activity, which these algorithms can leverage to differentiate between motor tasks more effectively.
| Subjects | Random For. (%) | SVM (%) | XGBoost (%) | LDA (%) | ||||
| EEG | tEEG | EEG | tEEG | EEG | tEEG | EEG | tEEG | |
| S1 | 65 | 82.22 | 73.33 | 76.67 | 63.33 | 78.89 | 40 | 67.78 |
| S2 | 64.1 | 87.47 | 67.18 | 71.79 | 61.03 | 84.32 | 48.21 | 72.74 |
| S3 | 73.86 | 89.37 | 78.3 | 79.6 | 71.57 | 87.92 | 60.26 | 76.47 |
| S4 | 90.83 | 95.56 | 90 | 80.56 | 87.5 | 94.44 | 64.17 | 66.11 |
| S5 | 72.92 | 86.94 | 72.5 | 65.56 | 71.25 | 85.83 | 62.5 | 66.39 |
| S6 | 78 | 89.33 | 69 | 76 | 75 | 88.67 | 61 | 69.33 |
| S7 | 91.88 | 97.92 | 88.75 | 85.83 | 90 | 97.08 | 80 | 77.08 |
| S8 | 69.38 | 85.83 | 73.12 | 64.17 | 75 | 85.83 | 53.75 | 71.25 |
| S9 | 69.38 | 82.92 | 70.62 | 72.92 | 66.25 | 82.08 | 54.37 | 64.17 |
| S10 | 70.62 | 81.25 | 72.5 | 69.17 | 65 | 80.83 | 47.5 | 64.58 |
| Mean | 74.60 | 87.88 | 75.53 | 74.23 | 72.59 | 86.59 | 57.18 | 69.59 |
@@
Table 4 summarizes the classification accuracies for a multiclass scenario involving No-movement versus different grasp tasks, contrasting the performance between EEG and tEEG across various subjects and machine learning models. It is notable that almost all classification results exceed the chance level of 33.33%, indicating that both EEG and tEEG provide signals containing sufficient information to distinguish between the classes above a level of random guessing. This is an important baseline as it establishes that the neural signals being classified carry task-relevant information. tEEG consistently shows an advantage over conventional EEG across all subjects and models, with the mean accuracy for tEEG being notably higher. The Random Forest classifier in particular reveals a substantial mean accuracy increase from 47.06% with conventional EEG to 72.51% with tEEG. The performance of the SVM model exhibits robustness across both EEG and tEEG modalities, with nearly identical average results, although EEG displays a marginally higher accuracy. XGBoost and LDA models further reinforce this trend, showing substantial improvements when using tEEG data, with mean accuracies of 75.97% and 63.46% respectively.
| Subjects | Random For. (%) | SVM (%) | XGBoost (%) | LDA (%) | ||||
| EEG | tEEG | EEG | tEEG | EEG | tEEG | EEG | tEEG | |
| S1 | 38.67 | 63.7 | 57.33 | 51.85 | 41.33 | 79.26 | 20 | 72.59 |
| S2 | 36.25 | 70.91 | 56.25 | 58.28 | 47.5 | 76.31 | 33.75 | 63.94 |
| S3 | 47.27 | 76.78 | 60.91 | 68.12 | 55.45 | 76.29 | 37.27 | 71.77 |
| S4 | 63.33 | 75.19 | 72 | 67.04 | 68 | 87.04 | 43.33 | 75.56 |
| S5 | 42 | 68.89 | 61 | 57.22 | 46.33 | 65 | 46.33 | 50.56 |
| S6 | 53.6 | 72.44 | 59.2 | 63.11 | 58.4 | 79.11 | 35.2 | 61.33 |
| S7 | 63 | 84.72 | 68.5 | 73.89 | 66.5 | 82.22 | 69 | 70.83 |
| S8 | 42.5 | 71.67 | 60 | 49.17 | 48 | 73.89 | 39.5 | 60.56 |
| S9 | 40 | 70 | 60 | 58.06 | 43.5 | 73.61 | 46 | 53.61 |
| S10 | 44 | 70.83 | 57.5 | 54.44 | 46 | 66.94 | 38 | 53.89 |
| Mean | 47.06 | 72.51 | 61.27 | 60.12 | 52.10 | 75.97 | 40.84 | 63.46 |
@@
The bar plot in Figure 7 illustrates the improvement in classification accuracy achieved by using tEEG over conventional EEG for various binary and multiclass tasks. Each bar represents the percentage increase in accuracy across different machine learning models. The data suggest that while SVM is relatively robust, with smaller improvements seen with tEEG in some scenarios, the other models exhibit substantial accuracy enhancements when employing tEEG. Specifically, the improvement is more pronounced for RF, XGBoost, and LDA models. Furthermore, the bar plot clearly indicates that the superiority of tEEG is more pronounced in complex classification tasks. In the context of multiclass classification, where the task difficulty is inherently higher due to the increase in the number of classes, the advantages of tEEG become more evident. For instance, the RF model shows a remarkable improvement, which is the highest across all tasks and models represented.

### v2

@@ 2.5 Feature Extraction and Classification
The primary objective during the feature extraction phase was to capture relevant information from the EEG and tEEG data. The time window [-1, 1] s was defined as the time region of interest (tROI), with 0 s corresponding to the movement onset. Wavelets, specifically the complex Morlet wavelet (cmor), were employed to decompose these segments into various frequency bands: delta, theta, alpha, beta, and gamma. For each of the identified frequency bands in every channel, four statistical features were extracted from the magnitude of the wavelet coefficients. These features are: mean of the magnitude, variance of the magnitude, skewness of the magnitude, and kurtosis of the magnitude.
Given that there were 4 channels, each channel containing data from 5 frequency bands, and each band producing 4 features, the total feature count amounted to 80 for every data instance. With the features in hand, the subsequent step was classification. We employed four distinct machine learning algorithms: Support Vector Machine (SVM), Random Forest (RF), Extreme Gradient Boosting (XGBoost), and Linear Discriminant Analysis (LDA). These techniques were used for both multiclass and binary classification scenarios. To assess the effectiveness of these models, we implemented a 5-fold cross-validation method. This approach helped us gain a reliable measure of the models’ ability to generalize, as it calculated the average performance indicators across all the validation folds. For all comparisons and analyses in this study, paired t-tests were employed to statistically validate the results.
@@ 3.1 Comparative Analysis
The comparative analysis of SNRs between the tEEG and conventional EEG technologies revealed a consistent pattern favoring the tEEG method across various channels. As delineated in Figure 3, the grand average SNR values for all subjects besides the p-values were assessed at channels P3, C3, C4, and F5. Notably, the tEEG (green boxes) significantly exhibited higher SNRs across all channels when compared to the conventional EEG (blue boxes). Specifically, the median SNR for the tEEG ranged from 56.24 dB at channel C3 to 59.56 dB at channel P3, which was the highest observed across all channels and technologies. In comparison, the conventional EEG technology presented lower median SNRs, with the lowest at channel C4 (53.92 dB) and the highest at channel P3 (54.57 dB). In addition to the improved median SNRs, the tEEG also demonstrated lower variability in SNR measurements among subjects, as evidenced by the narrower interquartile ranges in the green boxes. This reduced variance suggests that the tEEG provides more consistent and reliable SNR values, which is a critical aspect of EEG signal quality.
Subjects Random For. (%) SVM (%) XGBoost (%) LDA (%) EEG tEEG EEG tEEG EEG tEEG EEG tEEG S1 70.00 90.00 75.00 75.56 73.33 88.89 51.67 78.89 S2 82.82 87.42 76.54 85.42 84.36 84.32 56.28 72.89 S3 71.76 92.36 74.05 83.3 76.34 90.17 50.26 84.10 S4 95.00 96.11 89.17 85.56 90.83 97.22 68.33 76.67 S5 69.58 90.56 69.17 68.06 67.08 86.11 60.42 69.44 S6 80.00 91.33 73.00 78.67 76.00 88.67 43.00 78.67 S7 94.38 98.24 93.12 96.67 92.50 98.33 84.38 89.58 S8 73.75 89.58 70.00 70.00 71.25 90.00 66.88 73.75 S9 73.12 91.67 67.50 72.92 65.62 90.00 59.38 72.08 S10 68.12 87.08 66.88 75.83 68.12 82.50 60.00 67.50 Mean
Subjects Random For. (%) SVM (%) XGBoost (%) LDA (%) EEG tEEG EEG tEEG EEG tEEG EEG tEEG S1 76.67 90.00 78.33 85.56 78.33 90.00 51.67 71.11 S2 78.08 88.53 62.56 82.21 67.18 84.37 53.21 76.11 S3 75.03 93.22 62.35 82.56 66.99 93.19 48.95 71.20 S4 93.33 92.78 88.33 82.22 91.67 92.22 64.17 77.78 S5 70.42 88.89 72.08 76.94 69.17 84.72 63.33 69.72 S6 80.00 88.67 67.00 87.33 75.00 90.67 62.00 76.00 S7 93.75 99.17 92.50 94.58 93.12 97.92 80.62 92.92 S8 68.12 89.58 70.62 73.33 64.38 88.75 54.37 67.50 S9 70.62 89.17 68.12 77.08 67.50 88.75 65.00 70.42 S10 71.25 87.08 69.38 77.92 67.50 86.67 54.37 77.08 Mean 77.73 8.67 90.71 3.34 73.13 9.72 81.97 5.84 74.08 9.97 89.73 3.89 59.77 8.85 74.98 6.86
Subjects Random For. (%) SVM (%) XGBoost (%) LDA (%) EEG tEEG EEG tEEG EEG tEEG EEG tEEG S1 65.00 82.22 73.33 76.67 63.33 78.89 40.00 67.78 S2 64.10 87.47 67.18 71.79 61.03 84.32 48.21 72.74 S3 73.86 89.37 78.30 79.60 71.57 87.92 60.26 76.47 S4 90.83 95.56 90.00 80.56 87.50 94.44 64.17 66.11 S5 72.92 86.94 72.50 65.56 71.25 85.83 62.50 66.39 S6 78.00 89.33 69.00 76.00 75.00 88.67 61.00 69.33 S7 91.88 97.92 88.75 85.83 90.00 97.08 80.00 77.08 S8 69.38 85.83 73.12 64.17 75.00 85.83 53.75 71.25 S9 69.38 82.92 70.62 72.92 66.25 82.08 54.37 64.17 S10 70.62 81.25 72.50 69.17 65.00 80.83 47.50 64.58 Mean 74.60 9.22 87.88 5.20 75.53 7.47 74.23 6.5 72.59 9.26 86.59 4.45 57.18 10.54 69.59 4.43
Subjects Random For. (%) SVM (%) XGBoost (%) LDA (%) EEG tEEG EEG tEEG EEG tEEG EEG tEEG S1 38.67 63.70 57.33 51.85 41.33 79.26 20.00 72.59 S2 36.25 70.91 56.25 58.28 47.50 76.31 33.75 63.94 S3 47.27 76.78 60.91 68.12 55.45 76.29 37.27 71.77 S4 63.33 75.19 72.00 67.04 68.00 87.04 43.33 75.56 S5 42.00 68.89 61.00 57.22 46.33 65.00 46.33 50.56 S6 53.60 72.44 59.20 63.11 58.40 79.11 35.20 61.33 S7 63.00 84.72 68.50 73.89 66.50 82.22 69.00 70.83 S8 42.50 71.67 60.00 49.17 48.00 73.89 39.50 60.56 S9 40.00 70.00 60.00 58.06 43.50 73.61 46.00 53.61 S10 44.00 70.83 57.50 54.44 46.00 66.94 38.00 53.89 Mean 47.06 9.24 72.51 5.27 61.27 4.80 60.12 7.40 52.10 9.00 75.97 6.28 40.84 11.81 63.46 8.49
@@
Figure 4 presents the grand average functional connectivity between EEG channels, quantified across all subjects for both EEG and tEEG technologies. The heatmaps visually compare the connectivity strengths, with conventional EEG displayed on the top and tEEG on the bottom. The conventional EEG heatmap shows considerable functional connectivity, with several channel pairs exhibiting strong correlations, such as P3-F5 (0.82) and C3-F5 (0.8), suggesting significant shared signal components across these channels. This is indicative of a lower spatial resolution, where the activity from different brain regions may be detected simultaneously by multiple channels. Conversely, the tEEG heatmap illustrates substantially lower connectivity values for all channel pairs when compared to conventional EEG. This is particularly evident in the markedly reduced correlations for channel pairs P3-F5 (0.01) and C3-F5 (0.048).
@@
Figure 5 showcases the wavelet time-frequency representations for conventional EEG and tEEG, analyzing the neural dynamics from one second before to one second after the onset of movement, which is marked by time zero on the x-axis. During the Power Grasp and Precision Grasp conditions, the time-frequency maps exhibit a high increase in activity, particularly noticeable around 500ms post-movement onset. The tEEG maps display a more significant increase in activity across a range of frequencies when compared to conventional EEG. This increased activity, especially evident in the beta frequency band, is indicative of the sensorimotor processing associated with the execution of the grasping tasks. In the No Movement condition, conventional EEG exhibits minimal activity, aligning with expectations of neural quiescence. However, tEEG reveals subtle low-frequency activities even in this resting state, indicating its heightened sensitivity to neural oscillations that conventional EEG may not capture.
@@ 3.2 Feature Extraction and Classification
Table 1 presents the binary classification accuracies for distinguishing between No-movement and Power Grasp tasks across different subjects using various machine learning models with accuracies reported for both conventional EEG and tEEG. Across all subjects, tEEG consistently outperformed conventional EEG in classification accuracy. The mean accuracy improvements with tEEG ranged from 3.76% in SVM to a notable 16.30% in LDA, indicating a substantial enhancement in the machine learning models’ ability to utilize the data captured by tEEG. The Random Forest classifier achieved the highest mean accuracy for tEEG at 91.53%. Similarly, LDA showed the least improvement, yet still produced a mean accuracy increase of 16.30% with tEEG. Individual subject analysis revealed that Subject 7 (S7) exhibited the highest classification accuracy with tEEG at 98.33% using the XGBoost model, while the lowest was for Subject 10 (S10) at 67.5% using LDA.
@@
Table 2 details the binary classification accuracies for No-movement versus Precision Grasp tasks, utilizing a range of machine learning models and comparing the performance between conventional EEG and tEEG. The data illustrate a consistent trend where tEEG provides a marked improvement in classification accuracy over conventional EEG. The mean accuracies are significantly higher with tEEG, exhibiting an increase of 12.98% for Random Forest, 8.84% for SVM, 15.65% for XGBoost, and 15.21% for LDA. The XGBoost classifier shows the most substantial average increase in accuracy when using tEEG. Notably, every subject experienced an accuracy improvement with tEEG, affirming its universal enhancement of classification performance. Subject 7’s results stand out with an exceptional 99.17% accuracy using the Random Forest classifier in conjunction with tEEG, while the lowest accuracy with this technology is seen in Subject 8, with 67.50% using LDA. This spread suggests that while tEEG generally improves classification outcomes, the extent of its impact can vary among individuals.
@@
Table 3 summarizes the binary classification accuracies for discriminating between Power Grasp and Precision Grasp tasks, utilizing various machine learning models for both conventional EEG and tEEG. A key observation from the table is that while tEEG demonstrates an overall improvement in classification accuracy for most models, the SVM model with conventional EEG data exhibits a slightly higher average accuracy (75.53%) compared to tEEG (74.23%). However, this trend does not hold across other models. The XGBoost classifier, in particular, shows a marked increase in mean accuracy from 72.59% with EEG to 86.59% with tEEG. Subject 7’s classification with tEEG using Random Forest underscores this potential. For Random Forest and LDA models, tEEG also outperforms conventional EEG, with mean accuracies higher by 13.28% and 12.41%, respectively.
@@
Table 4 summarizes the classification accuracies for a multiclass scenario involving No-movement versus different grasp tasks, contrasting the performance between EEG and tEEG across various subjects and machine learning models. It is notable that almost all classification results exceed the chance level of 33.33%, indicating that both EEG and tEEG provide signals containing sufficient information to distinguish between the classes above a level of random guessing. This is an important baseline as it establishes that the neural signals being classified carry task-relevant information. tEEG consistently shows an advantage over conventional EEG across all subjects and models, with the mean accuracy for tEEG being significantly higher. The Random Forest classifier in particular reveals a substantial mean accuracy increase from 47.06% with conventional EEG to 72.51% with tEEG. The performance of the SVM model exhibits robustness across both EEG and tEEG modalities, with nearly identical average results, although EEG displays a marginally higher accuracy. XGBoost and LDA models further reinforce this trend, showing substantial improvements when using tEEG data, with mean accuracies of 75.97% and 63.46% respectively.
@@
Figure 6 illustrates the improvement in classification accuracy achieved by employing tEEG over conventional EEG for various binary and multiclass tasks. The corresponding p-values obtained from statistical analysis using a paired t-test are also presented, indicating the statistical significance of the observed performance differences. Each bar represents the percentage increase in accuracy across different machine learning models. The data suggest that while SVM is relatively robust, with smaller improvements seen with tEEG in some scenarios, the other models exhibit significantly accuracy enhancements when employing tEEG. Specifically, the improvement is more pronounced for RF, XGBoost, and LDA models.